The length of a rectangular mural is 2 meters greater than the width. The area is 48 square meters. Find the dimensions of the mural.
Answer by Boreal(15235)

length is x
width is x-2
area is product x^2-2x=48 and units are m*m=m^2
x^2-2x-48=0
(x-8)(x+6)=0
x=8,-6, and only 8 meters works for length
6 meters for width
Their product is 48 m^2.